in the javelin throw, hammer throw or discus the goal is maximum distance ignoring air resistence and the athlete's biomechanics at whatangle should the athlete throw these objects to maximize the distance they travel before hitting the ground
1
Expert's answer
2020-12-10T11:09:21-0500

The range is given by the formula


"R=\frac{v^2\sin{2\theta}}{g}"

Maximum distance will be when


"\sin{2\theta}=1\\\theta=45\degree"